OpenStudy (anonymous):

I have the answer can someone check to see if I;m right Which of the following is f ' (x) for f(x)=-3x^-5

OpenStudy (whpalmer4):

What's your answer?

OpenStudy (anonymous):

-15x^-4

OpenStudy (whpalmer4):

Sorry, that's incorrect.

OpenStudy (whpalmer4):

\[\frac{d}{dx}[u^n] = nu^{n-1}\]

OpenStudy (anonymous):

Ok so the is the answer 12x^-4

OpenStudy (whpalmer4):

No. What does n = ?

OpenStudy (anonymous):

n=-3

OpenStudy (whpalmer4):

No, isn't it \[f(x) = -3x^{-5}\]?

OpenStudy (anonymous):

yes so then doesn=-5

OpenStudy (whpalmer4):

Yes!

OpenStudy (whpalmer4):

Don't worry about the -3 in front for the moment, what is the derivative of \(x^{-5}\)?

OpenStudy (anonymous):

-4

OpenStudy (whpalmer4):

No...it's going to be some function of x

OpenStudy (anonymous):

Like-6

OpenStudy (whpalmer4):

No...it's not just a number, it needs to have an x in there!

OpenStudy (anonymous):

oh ok I gotha so like -8x

OpenStudy (whpalmer4):

No.... \[\frac{d}{dx}[x^n] = nx^{n-1}\]Your value of n is n = -5, right? \[\frac{d}{dx}[x^{-5} = -5x^{-5-1} = 5x^{-6}\]

OpenStudy (anonymous):

ok so then what or is the 5x^-6 my answer

OpenStudy (whpalmer4):

Nuts, missed a bracket: \[\frac{d}{dx}[x^{-5}] = -5x^{-5-1} = -5x^{-6}\] So what does \[\frac{d}{dx}[-3x^{-5}] = \]?

OpenStudy (anonymous):

ok so I put it into my calculator bt for the space in the calculator where it says x= at the end what do I put

OpenStudy (whpalmer4):

You don't need a calculator for this... If the derivative of x^-5 = -5x^-6, what is the derivative of -3x^-5?

OpenStudy (anonymous):

This is where I get confused can you explain this to me

OpenStudy (whpalmer4):

Remember, the definition of the derivative is \[\lim_{h\rightarrow0}\frac{f(x+h)-f(x)}{h}\] Let's say our f(x) is simply f(x) = x. Then by the definition, we have: \[\lim_{h\rightarrow0}\frac{f(x+h)-f(x)}{h} =\lim_{h\rightarrow0}\frac{(x+h)-(x)}{h} = 1 \] right?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Now let's say our f(x) is f(x) = 3x. \[\lim_{h\rightarrow0}\frac{f(x+h)-f(x)}{h} =\lim_{h\rightarrow0}\frac{3(x+h)-3(x)}{h} =\]\[3 \lim_{h\rightarrow0}\frac{(x+h)-(x)}{h} = 3*1 = 3 \] right?

OpenStudy (anonymous):

so then would the answer be -15x^-6

OpenStudy (whpalmer4):

what is -3 * -5?

OpenStudy (whpalmer4):

watch out forehead, here comes palm :-)

OpenStudy (anonymous):

positive 15x^-6

OpenStudy (whpalmer4):

yes!

OpenStudy (anonymous):

ok thanks

OpenStudy (whpalmer4):

So I wasn't sure if your initial mistake was subtracting 1 from the leading -3 instead of the exponent, or if you were adding 1 instead of subtracting one from the -5...

OpenStudy (anonymous):

I think it was where you had o subtract 1 from -3

OpenStudy (whpalmer4):

Here's another one: if f(x) = 3x^2, what is f'(x) =?

OpenStudy (anonymous):

would it be 6x

OpenStudy (whpalmer4):

correct again! One more: f(x) = -1x^-1, f'(x) = ?

OpenStudy (anonymous):

1x

OpenStudy (whpalmer4):

Nope. Try again.

OpenStudy (anonymous):

x^-2

OpenStudy (whpalmer4):

Yes. Remember, the exponent always decreases.

OpenStudy (anonymous):

ok but I have another problem but with a fraction can you help me with that one

OpenStudy (whpalmer4):

Sure, as long as it is a quick one...

OpenStudy (anonymous):

yes ok same instructions but f '(x) for f(x)=(1/x^7)

OpenStudy (whpalmer4):

\[f(x) = \frac{1}{x^7}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the answerx^-7

OpenStudy (whpalmer4):

Remember the property of exponents: \[a^{-n} = \frac{1}{a^n}\]

OpenStudy (anonymous):

ok so then the answer would be x^-7

OpenStudy (whpalmer4):

How can you rewrite your problem given that?

OpenStudy (anonymous):

what I;m confused

OpenStudy (whpalmer4):

\[f(x) = \frac{1}{x^7} = x^{-7}\]

OpenStudy (whpalmer4):

Now take the derivative

OpenStudy (anonymous):

-7x^6

OpenStudy (whpalmer4):

Okay, you're not following the rules: \[\frac{d}{dx}[u^n] = nu^{n-1}\]

OpenStudy (anonymous):

(1/-7x^6

OpenStudy (whpalmer4):

I should be writing that with a du, not a dx, but whatever... If your initial exponent is n = -7, then the result is \(-7x^{-7-1} = -7x^{-8}\)

OpenStudy (anonymous):

oh ok for some reason I was doing -7+1 and that was where I was getting the -6 from

OpenStudy (whpalmer4):

Yes, I figured as much. So, don't do that :-) If your answer is \(-7x^{-8}\) what would that be written as a fraction?

OpenStudy (anonymous):

(-7/x^8)

OpenStudy (anonymous):

is that correct

OpenStudy (whpalmer4):

Yes. I was hoping you wouldn't make the mistake of writing 1/(-7x^8) :-)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

thanks

OpenStudy (whpalmer4):

You're welcome. Remember, watch the subtraction from the exponent carefully!

OpenStudy (anonymous):

ok thnks